OpenStudy (pphalke):

find the x and y intercept for the following equation and write an ordered pair for each, round if necessary 2x+3y=6

Directrix (directrix):

The x-intercept is where the line crosses the x-axis. To get on the x-axis, the y-value of the x-intercept must be 0. 2x+3y=6 x-intercept Let y = 0 2x + 3*0 =6 Solve this for x.

OpenStudy (pphalke):

2x+0=6?

OpenStudy (pphalke):

then you subtract 0 from both sides?

OpenStudy (pphalke):

then 2x=6 then x=3

OpenStudy (pphalke):

@Directrix

OpenStudy (hyperpiper):

You are correct :)

Directrix (directrix):

The x-intercept has coordinates (3,0)

OpenStudy (pphalke):

how do i find y intercept

OpenStudy (hyperpiper):

plug in 0 for x

Directrix (directrix):

The y-intercept is on the y-axis. To get on the y-axis, the x-coordinate must be 0.

OpenStudy (hyperpiper):

and you do the same thing as in the other equation

Directrix (directrix):

Solve for y: 2x+3y=6 Let x = 0 2*0 + 3y = 6 y = ?

OpenStudy (pphalke):

0+3y=6 what do i do from there do i sub 6 or 3 from both sides

Directrix (directrix):

3y = 6 y = ?

OpenStudy (pphalke):

y=2 so y int (2,0)

Directrix (directrix):

Coordinates of y-intercept: (0,2)

Directrix (directrix):

>>so y int (2,0) No. That point is on the x-axis. The y-intercept is on the y-axis and has x-coordinate 0.

OpenStudy (hyperpiper):

because (x,y) rules :)

OpenStudy (pphalke):

yes and thanks its right answer